a stone is thrown upwards with an initial velocity of 25m/s at an angle of 30 degree to the ground
how long will it take betweeeen when the stone was thrown and when it comes back to the ground?




1
Expert's answer
2022-02-02T14:38:08-0500

The time of flight is


"t=\dfrac{2v\sin\theta}g=2.55\text{ s}."